ON THE DISTRIBUTION OF THE EXTREMES OF UNEQUALLY CORRELATED NORMAL VARIABLES WITH APPLICATIONS TO ANTEDEPENDENT CLUSTER DATA

KRISHNA K. SAHA AND B. C. SUTRADHAR

Department of Mathematics and Statistics, Memorial University of Newfoundland, St. John’s, Newfoundland, Canada A1C 5S7

(Received December 16, 1996; revised August 21, 1997)

Abstract. Based on a small correlations approach, this paper deals with the probability distributions of the order statistics under a general Gaussian model with unequal (positive or negative) correlations. Illustrations are given for the cluster data those follow suitable antedependence models.

Key words and phrases: Unequal but small correlations, distribution of order statistics, tail probability.